OpenStudy (anonymous):

4y = 20 + 2y help please

OpenStudy (anonymous):

Y= 10

OpenStudy (ageta):

y=10

OpenStudy (ageta):

Let's solve your equation step-by-step. 4y=20+2y Step 1: Simplify both sides of the equation. 4y=2y+20 Step 2: Subtract 2y from both sides. 4y−2y=2y+20−2y 2y=20 Step 3: Divide both sides by 2.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

This website always helps me with this kind of stuff http://www.mathpapa.com/algebra-calculator.html

OpenStudy (ageta):

your welcome

OpenStudy (anonymous):

Helplpful website:- http://www.acalculator.com/logarithm-calculator-logx-logarithmic-equations.html